# Quantum-enhanced sensing of spin-orbit coupling without fine-tuning

### v2

@@ 4.1 Single particle probe
Figure 1: Scaling analysis of energy gap and QFI. Probe state is the ground state of the Hamiltonian in Eqs. (6)-(9) for sensing $\alpha_{z}$ with hopping parameter $t{=}1$, and $\alpha_{y}{=}\alpha_{z}$. (a) Energy gap between first excited state and ground state $\Delta$ with $\alpha_{z}$ and $B$ for system size $L{=}100$. The four points in the parameter space are chosen as $(B,\alpha_{z})=(0.05,0.1),(0.05,0.9),(0.9,0.9),(0.9,0.05)$, denoted by the cross, diamond, plus, and circle, respectively. (b) Scaling relation of the energy gap with system size at these four parameter points, labeled by corresponding legends. In all cases numerical fit $\Delta{=}aL^{-\mu}{+}b$ shows almost quadratic scaling of the gap closing. (c) QFI of the probe with respect to $\alpha_{z}$ in the parameter space spanned by $\alpha_{z}$ and $B$ for system size $L{=}100$. The same four points are chosen to study the scaling behavior. (d) Numerical fit confirms the quadratic scaling of QFI in the whole range as $F^{Q}{\sim}L^{\beta}$ with $\beta{\approx}2$.
@@
We start the sensing analysis with the single particle ground states of the Hamiltonian in Eq.(6) as the probe. In Fig. 1, we show the behavior of the energy gap $\Delta$ and QFI $F^{Q}$ of the ground state with respect to $\alpha_{z}$. We have considered experimentally relevant values of $\alpha_{z}$ and $B$ for our analysis. Fig 1(a) shows the result for a fixed size $L=100$ where, for weak magnetic fields, the gap goes through a minima at a small value of $\alpha_{z}$. As the field becomes stronger, the minima becomes less pronounced and the gap increases almost monotonically with $\alpha_{z}$. Interestingly, the scaling of the decline of energy gap with system size shows similar behavior across the entire range of $\alpha_{z}$ and $B$. This is shown in Fig. 1(b), where four different combinations of $\alpha_{z}$ and $B$ values in the weak and strong regimes have been chosen to analyze the scaling. In all the cases, the gap closing was found to scale almost quadratically with system size, i.e. $\Delta\sim L^{-\mu}$ with $\mu\approx 2$. Now for the QFI, Fig. 1(c) shows almost a complementary picture to Fig. 1(a), illustrating the close connection between the gap closing and the QFI. This leads to a quadratic scaling of QFI in the whole region spanned by $\alpha_{z}$ and $B$, i.e. $F^{Q}\sim L^{\beta}$ with $\beta\approx 2$, see Fig. 1(d). This is a very promising result as unlike criticality-based quantum sensors, where quantum-enhanced sensitivity is only achievable at the vicinity of the phase transition, one can reach Heisenberg scaling over a wide range of parameters. This can be explained by the favorable gap closing nature of the system. We also note that for a given system size, the sensing capacity is stronger for smaller values of SOC parameters in the presence of a weak external magnetic field.
4.2 Many-body interacting probe
@@
$H_{\rm int}=U\sum_{j}c^{{\dagger}}_{j,\uparrow}c_{j,\uparrow}+V\sum_{j,\sigma,\sigma^{\prime}}c^{{\dagger}}_{j,\sigma}c_{j,\sigma}c^{{\dagger}}_{j+1,\sigma^{\prime}}c_{j+1,\sigma^{\prime}}.$ (11)
The $U$ term accounts for an increase in the system’s energy when two fermions with opposite spins occupy the same site. For fermions in optical lattices, this term can be tuned by changing the scattering length with Feshbach resonance [87]. The $V$ term can arise due to the Coulomb repulsion, for example, between electrons confined in quantum dot arrays. As a typical scenario in theoretical and experimental situations, we consider the half-filled case here, i.e. $L$ fermions on a $L$-site wire. When $U{=}V{=}0$, the fermionic ground state is given by the antisymmetric Slater determinant state formed by the single particle energy eigenstates {|$\{\ket{E_{n}}\}$, which can be compactly written with the symmetric group $S_{L}$ as
|$\ket{\psi_{\rm GS}}=\frac{1}{\sqrt{L!}}\sum_{\bm{n}\in{S_{L}}}\text{sgn}(\bm{n})\ket{E_{n_{1}}}\dots\ket{E_{n_{L}}}.$ (12)
This results in an analytical QFI expression [78],
$F^{Q}=4\left(\sum_{l=1}^{L}\braket{\partial E_{l}}{\partial E_{l}}-\sum_{l,l^{\prime}}\braket{\partial E_{l}}{E_{l^{\prime}}}\braket{E_{l^{\prime}}}{\partial E_{l}}\right).$ (13)
@@
Figure 2: Many-body probe. Probe state is the half-filled ground state of the Hamiltonian in Eq. (10) with $t{=}1$, $B{=}0.01t$, and $\alpha_{y}{=}\alpha_{z}$. (a) Scaling of energy gap in the non-interacting case ($U{=}V{=}0$) with system size $L$ for different values of SOC parameter $\alpha_{z}$. (b) Scaling of QFI with respect to $\alpha_{z}$ in the non-interacting case for different $\alpha_{z}$ values. (c) QFI at $\alpha_{z}{=}0.1t$ as a function of $U$ when $V{=}0$. (d) QFI scaling at $\alpha_{z}{=}0.1t$ for different $U$ values. (e) Spectrum $E$ as a function of $U$. (f) QFI at $\alpha_{z}{=}0.1t$ as a function of $V$ when $U{=}0$.
@@
Top row of Fig. 2 shows the results of the many-body non-interacting probe where the external field is fixed at $B{=}0.01t$. As the single particle eigenstates are filled up to $L$-th level in the ground state, the energy gap is defined by the energy difference between the $L{+}1$ and $L$-th eigenvalues. As Fig. 2(a) shows, gap closing is still algebraic with system size $(\Delta{\sim}L^{-\mu})$, but it is less than quadratic. The exponent $\mu$ is found to be decreasing with increasing $\alpha_{z}$. On the other hand, the QFI still scales quadratically with system size, as shown in Fig. 2(b). This shows that, the fermionic statistics in the non-interacting many-body ground state still results in the Heisenberg limit of precision for a wide range of the SOC parameter. When only $U$ is non-zero, we observe that the QFI starts increasing with $U$ before attaining a saturation in the repulsive case. For attractive $U$, the QFI decreases and goes to zero. Fig 2(c) displays this behavior, while Fig. 2(d) shows that algebraic scaling of QFI is still sustained in the repulsive case. The exponential growth of the Hilbert space dimension limits our calculations in terms of system size, which can therefore be affected by finite size effects. Nevertheless, the quadratic scaling of QFI for $U{=}0$ is expectedly observed in Fig. 2(d). Surprisingly, as $U$ is increased, the decrease in exponent $\beta$ is not drastic, which indicates that the interacting probes can still provide quantum enhancement. This can be attributed to the fact that the presence of $U$ in Eq. (11) does not open up the ground state energy gap as shown in the band structure in Fig. 2(e). The fermions in the ground state stay localized at different sites. In the other case, when $V$ is non-zero and $U{=}0$, the QFI steadily decreases (see Fig. 2(f)). We numerically observe that quadratic scaling of QFI survives for small values of $V$ and disappears near $V{\approx}t$.
4.3 Thermal probe
@@
Figure 3: Thermal probe. Probe state is the thermal state of the Hamiltonian in Eq. (10) with $t{=}1$, $B{=}0.01t$, $\alpha_{y}{=}\alpha_{z}$. (a) QFI vs. temperature in the non-interacting case for different system sizes. As thermal energy surpasses the gap, QFI falls of universally as ${\sim}1/T$. (b) Variation of QFI with both $U$ and $T$ for $L{=}6$.
@@
So far we have discussed the cases where the probe is taken to be the ground state which is the equilibrium state at zero temperature. In practice however, the sensors generally operate at finite temperature and hence, it is necessary to analyze the effect of temperature on the sensing capability. When the temperature is smaller than the energy gap i.e. $k_{B}T{<}\Delta$ ($k_{B}$ is the Boltzmann constant), we expect the ground state description to hold. Higher temperatures result in a mixed state with increasing contributions from higher energy states. The QFI then starts to decrease and typically collapses universally with an algebraic decay with temperature [74]. In Fig. 3(a), we show this behavior in the non-interacting case for two system sizes $L=100,200$. The ground state QFI value is retained up to a temperature $T{\sim}\Delta/k_{B}$, after which the QFI is found to fall off as $T^{-1}$. In the repulsive interacting case, we observe that the energy gap decreases with $U$, and hence the temperature at which the QFI starts to fall decreases with $U$ (see Fig. 3(b)).
5 Multi-parameter sensing
@@
Figure 4: Multi-parameter sensing. Probe state is the single particle ground state of the Hamiltonian in Eq. (6) with $t{=}1$, and $B{=}0.01t$. (a) The $(\alpha_{z},\alpha_{z})$ component of the QFIM $\mathcal{F}^{Q}$ across a range of $\alpha_{z}$ and $\alpha_{z}$ for $L{=}100$. Four points chosen in the parameter space for scaling analysis are $(\alpha_{y},\alpha_{z})=(0.05,0.1),(0.9,0.05),(0.9,0.9),(0.05,0.9)$, denoted by the cross, diamond, plus, and circle, respectively. (b) Scaling of $(\mathcal{F}^{Q})_{\alpha_{z},\alpha_{z}}$ at the four points. (c) The $(\alpha_{y},\alpha_{y})$ component of $\mathcal{F}^{Q}$. (d) Scaling of $(\mathcal{F}^{Q})_{\alpha_{y},\alpha_{y}}$ at the four points. (e) The $(\alpha_{y},\alpha_{z})$ component of $\mathcal{F}^{Q}$. (d) Scaling of $(\mathcal{F}^{Q})_{\alpha_{y},\alpha_{z}}$ at the four points.
@@
In this section, we consider a more general scenario where we do not presume prior knowledge of the SOC parameters i.e. the Rashba fields in any direction. The goal is to estimate both $\alpha_{y}$ and $\alpha_{z}$ simultaneously, making the problem multi-parametric. To analyze the system’s sensitivity to changes in these parameters, we compute the multi-parameter QFIM $\mathcal{F}^{Q}$, which provides a comprehensive description of the ultimate precision limits for estimating multiple parameters concurrently. In Fig. 4(a) we shows the component $\mathcal{F}^{Q}_{\alpha_{z},\alpha_{z}}$ as a function of the two parameters. Fig 4(b) shows that $\mathcal{F}^{Q}_{\alpha_{z},\alpha_{z}}$ reaches super-linear scaling in a wide range of $\alpha_{z}$ and $\alpha_{y}$ and can also attain close to Heisenberg scaling. Similar observation also holds for the component $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{y}}$, as shown in Figs. 4(c) and (d). From this analysis we can also infer that the quantum enhancement of the sensing of $\alpha_{y}$ when $\alpha_{z}$ is known, can reach the Heisenberg limit as well. In general, however, the sensitivity depends on the off-diagonal component $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{z}}$ that captures the correlation between the estimates of $\alpha_{z}$ and $\alpha_{y}$. Fig 4(e) and (f) show that the scaling of $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{z}}$ also stays close to quadratic across the whole range.
6 Optimal measurement

### v3

@@ IV.1 Single particle probe
Figure 1: Scaling analysis of energy gap and QFI. Probe state is the ground state of the Hamiltonian in Eqs. (7)-(10) for sensing $\alpha_{z}$ with hopping parameter t=1t{=}1, and $\alpha_{y}{=}\alpha_{z}$. (a) Energy gap between first excited state and ground state $\Delta$ with $\alpha_{z}$ and $B$ for system size L=100L{=}100. The four points in the parameter space are chosen as $(B,\alpha_{z})=(0.05,0.1),(0.9,0.1),(0.9,0.9),(0.05,0.9)$, denoted by the cross, diamond, plus, and circle, respectively. (b) Scaling relation of the energy gap with system size at these four parameter points, labeled by corresponding legends. In all cases numerical fit $\Delta{=}aL^{-\mu}{+}b$ shows almost quadratic scaling of the gap closing. (c) QFI of the probe with respect to $\alpha_{z}$ in the parameter space spanned by $\alpha_{z}$ and $B$ for system size L=100L{=}100. The same four points are chosen to study the scaling behavior. (d) Numerical fit confirms the quadratic scaling of QFI in the whole range as $F^{Q}{\sim}L^{\beta}$ with $\beta{\approx}2$.
@@
We start the sensing analysis with the single particle ground states of the Hamiltonian in Eq.(7) as the probe. In Fig. 1, we show the behavior of the energy gap $\Delta$ and QFI $F^{Q}$ of the ground state with respect to $\alpha_{z}$. We have considered experimentally relevant values of $\alpha_{z}$ and $B$ for our analysis. Fig. 1(a) shows the result for a fixed size $L=100$ where, for weak magnetic fields, the gap goes through a minimum at a small value of $\alpha_{z}$. As the field becomes stronger, the minimum becomes less pronounced and the gap increases almost monotonically with $\alpha_{z}$. Interestingly, the scaling of the decline of energy gap with system size shows similar behavior across the entire range of $\alpha_{z}$ and $B$. This is shown in Fig. 1(b), where four different combinations of $\alpha_{z}$ and $B$ values in the weak and strong regimes have been chosen to analyze the scaling. In all the cases, the gap closing was found to scale almost quadratically with system size, i.e. $\Delta\sim L^{-\mu}$ with $\mu\approx 2$. Indeed, for the QFI, Fig. 1(c) shows almost a complementary picture to Fig. 1(a), illustrating the close connection between the gap closing and the QFI. This leads to a quadratic scaling of QFI in the whole region spanned by $\alpha_{z}$ and $B$, i.e. $F^{Q}\sim L^{\beta}$ with $\beta\approx 2$, see Fig. 1(d). This is a very promising result as unlike criticality-based quantum sensors, where quantum-enhanced sensitivity is only achievable at the vicinity of the phase transition, one can reach Heisenberg scaling over a wide range of parameters. This can be explained by the favorable gap closing nature of the system. We also note that for a given system size, the sensing capacity is stronger for smaller values of SOC parameters in the presence of a weak external magnetic field. The measurement settings with which the error bounds given by the QFI can be approached is discussed later in Sec. VI.
IV.2 Many-body interacting probe
@@
$H_{\rm int}=U\sum_{j}c^{{\dagger}}_{j,\uparrow}c_{j,\downarrow}+V\sum_{j,\sigma,\sigma^{\prime}}c^{{\dagger}}_{j,\sigma}c_{j,\sigma}c^{{\dagger}}_{j+1,\sigma^{\prime}}c_{j+1,\sigma^{\prime}}.$ (12)
[…]
This results in an analytical QFI expression Sarkar et al. (2022),
$F^{Q}=4\left(\sum_{l=1}^{L}\braket{\partial E_{l}|\partial E_{l}}-\sum_{l,l^{\prime}}\braket{\partial E_{l}|E_{l^{\prime}}}\braket{E_{l^{\prime}}|\partial E_{l}}\right).$ (14)
@@
Figure 2: Many-body probe. Probe state is the half-filled ground state of the Hamiltonian in Eq. (11) with t=1t{=}1, B=0.01tB{=}0.01t, and $\alpha_{y}{=}\alpha_{z}$. (a) Scaling of energy gap in the non-interacting case (U=V=0U{=}V{=}0) with system size $L$ for different values of SOC parameter $\alpha_{z}$. (b) Scaling of QFI with respect to $\alpha_{z}$ in the non-interacting case for different $\alpha_{z}$ values. (c) QFI at $\alpha_{z}{=}0.1t$ as a function of $U$ when V=0V{=}0. (d) QFI scaling at $\alpha_{z}{=}0.1t$ for different $U$ values. (e) Spectrum $E$ as a function of $U$. (f) QFI at $\alpha_{z}{=}0.1t$ as a function of $V$ when U=0U{=}0.
@@
Top row of Fig. 2 shows the results of the many-body non-interacting probe where the external field is fixed at B=0.01tB{=}0.01t. As the single particle eigenstates are filled up to $L$-th level in the ground state, the energy gap is defined by the energy difference between the L+1L{+}1 and $L$-th eigenvalues. As Fig. 2(a) shows, gap closing is still algebraic with system size $(\Delta{\sim}L^{-\mu})$, but it is less than quadratic. The exponent $\mu$ is found to be decreasing with increasing $\alpha_{z}$. On the other hand, the QFI still scales quadratically with system size, as shown in Fig. 2(b). This shows that, the fermionic statistics in the non-interacting many-body ground state still results in the Heisenberg limit of precision for a wide range of the SOC parameter. When only $U$ is non-zero, we observe that the QFI starts increasing with $U$ before attaining a saturation in the repulsive case. For attractive $U$, the QFI decreases and goes to zero. Fig. 2(c) displays this behavior, while Fig. 2(d) shows that algebraic scaling of QFI is still sustained in the repulsive case. The exponential growth of the Hilbert space dimension limits our calculations in terms of system size, which can therefore be affected by finite size effects. Nevertheless, the quadratic scaling of QFI for U=0U{=}0 is expectedly observed in Fig. 2(d). Surprisingly, as $U$ is increased, the decrease in exponent $\beta$ is not drastic, which indicates that the interacting probes can still provide quantum enhancement. This can be attributed to the fact that the presence of $U$ in Eq. (12) does not open up the ground state energy gap as shown in the band structure in Fig. 2(e). The fermions in the ground state stay localized at different sites. In the other case, when $V$ is non-zero and U=0U{=}0, the QFI steadily decreases (see Fig. 2(f)). We numerically observe that quadratic scaling of QFI survives for small values of $V$ and disappears near $V{\approx}t$.
IV.3 Origin of quantum-enhancement
In order to understand the underlying mechanism for the quantum-enhanced sensing with ground state probes we rely on recent analytical results of Ref. Abiuso et al. (2025). According to these results, the QFI of the ground state of a Hamiltonian $H_{\alpha}{=}H_{1}{+}\alpha H_{2}$, is upper bounded by $H_{2}||^{2}/\Delta^{2}$, where $||{\bullet}||$ denotes the operator semi-norm (difference between highest and lowest eigenvalues) Abiuso et al. (2025). In our case, when considering the estimation of $\alpha_{y}{=}\alpha_{z}$, we have $H_{2}{=}\sum_{j,\sigma,\sigma^{\prime}}i(\sigma_{z}-\sigma_{y})_{\sigma,\sigma^{\prime}}c^{{\dagger}}_{j+1,\sigma}c_{j,\sigma^{\prime}}+\text{H.c.}$, and it can be checked quickly that $H_{2}||$ does not scale with system size. Therefore the enhancement solely comes form the scaling of the gap $\Delta{\sim}L^{-\mu}$ if $\mu{>}1/2$. While this only gives the upper bound, typically the same scaling behavior is displayed by the QFI near a phase transition Abiuso et al. (2025); Sarkar et al. (2025). Away from a transition, one needs to study on a case-by-case basis whether the quantum-enhancement is indeed manifested. For the single-particle probe, we observe that the energy gap falls quadratically with system size (Fig. 1(b)). Therefore, we expect the QFI to show a quantum-enhanced behavior as its upper bound scales as $L^{4}$. This shows that the observed quadratic scaling of QFI abides by the upper bound. Although the quadratic scaling of the QFI is sustained in a wide range of the SOC strangth, the scaling advantage is lost when SOC becomes the dominating term in the Hamiltonian. For example, when $\alpha_{z,y}\approx 10J$, the energy gap still falls as $L^{-2}$, but the ground state does not change appreciably with small changes in SOC strength. Therefore, the QFI is close to zero and becomes size-independent for strong SOC. For many-body non-interacting probe, the decrease of energy gap with system size occurs at a rate slower than quadratic fall-off (Fig. 2(a)). However, as the scaling is super-linear, we again expect the QFI to showcase the quantum enhancement. As shown in Fig. 2(b), we find that the QFI scales quadratically with system size, which is still within the upper bound.
Note that the resource with respect to which the scaling of the QFI is considered depends on the problem at hand. Subject to the context, this can be particle number, system size, or time. Although historically, Heisenberg limit has been obtained with respect to particle number Giovannetti et al. (2004), in a non-interacting ground state probe, it is possible to consider the system size instead as it determines the spatial extension over which the particle is delocalized. The system size and particle number can even be interchangeable in some scenarios such as a spin-system that can be mapped on to spin-less fermions on a lattice with Jordan-Wigner transformation. Although it is not applicable to our case, in a dynamics based probe, time serves as a resource for quantum enhancement Boixo et al. (2007); Ilias et al. (2022). For our model, it is possible to look additionally at the scaling with respect to particle numbers. However, it would lead to only linear scaling as the ground state would be a tensor product of fermionic single-particle eigenstates and the QFI is additive He et al. (2023); Manshouri et al. (2025).
IV.4 Thermal probe
@@
Figure 3: Thermal probe. Probe state is the thermal state of the Hamiltonian in Eq. (11) with t=1t{=}1, B=0.01tB{=}0.01t, $\alpha_{y}{=}\alpha_{z}$. (a) QFI vs. temperature in the non-interacting case for different system sizes. As thermal energy surpasses the gap, QFI falls of universally as ${\sim}1/T$. (b) Variation of QFI with both $U$ and $T$ for L=6L{=}6.
@@
So far we have discussed the cases where the probe is taken to be the ground state which is the equilibrium state at zero temperature. In practice however, the sensors generally operate at finite temperature and hence, it is necessary to analyze the effect of temperature on the sensing capability. When the temperature is smaller than the energy gap i.e. $k_{B}T{<}\Delta$ ($k_{B}$ is the Boltzmann constant), we expect the ground state description to hold. Higher temperatures result in a mixed state with increasing contributions from higher energy states. The QFI then starts to decrease and typically collapses universally with an algebraic decay with temperature He et al. (2023). In Fig. 3(a), we show this behavior in the non-interacting case for two system sizes $L=100,200$. The ground state QFI value is retained up to a temperature $T{\sim}\Delta/k_{B}$, after which the QFI is found to fall off as $T^{-1}$. In the repulsive interacting case, we observe that the energy gap decreases with $U$, and hence the temperature at which the QFI starts to fall decreases with $U$ (see Fig. 3(b)).
IV.5 Disorder
After analyzing the effect of thermal noise, we now consider another practical issue that can take place in the form of disorders during physical realization. The disorder can occur in various terms, including in the tunneling strength $t$ and in the applied magnetic field $B$, which is crucial for breaking the degeneracy. Moreover lattice defects can also take place in the form of holes or unoccupied sites. To see the effect of these disorders on the sensing, we include an additional term to the Hamiltonian in Eq. (7)
$H_{\rm dis}=-\sum_{j,\sigma}\Delta t_{j}(c^{{\dagger}}_{j+1,\sigma}c_{j,\sigma}+c^{{\dagger}}_{j,\sigma}c_{j+1,\sigma})$
$+\sum_{j,\sigma,\sigma^{\prime}}\Delta B_{j}\;(\sigma_{z})_{\sigma,\sigma^{\prime}}\;c^{{\dagger}}_{j,\sigma}c_{j,\sigma^{\prime}}+\sum_{j,\sigma}\mu_{j}c^{{\dagger}}_{j,\sigma}c_{j,\sigma}.$ (15)
Here $\Delta t_{j}$ is a random disturbance to the tunneling between sites $j$ and j+1j{+}1 that is drawn from a uniform distribution $0,W_{t}]$, with $W_{t}$ as the disorder strength. $\Delta B_{j}$ is a random magnetic field along $z$-direction at the $j$-th site that is drawn from a uniform distribution $W_{B},W_{B}]$, with $W_{B}$ as the disorder strength. Finally, $\mu_{j}$ is an onsite potential that takes a high value $\mu$ on some randomly chosen sites, thus making those sites practically unoccupied due to high energy penalties. We examine the effect of each type of disorder individually with $W_{t}{=}0.01t$, $W_{B}{=}0.1B$, and $\mu{=}10t$ with number density of holes as $L/50$. We find these values to be typical disorder strengths and hole density up to which quantum enhancement is observed. The departure from quadratic scaling due to disorder is shown in Fig. 4 for the single particle probe. The four panels Figs. 4(a), (b), (c), (d) correspond to the four points in the $B-\alpha_{z}$ plane that have been considered before as well. In each panel, the scaling behavior is shown for the individual cases of disorder in $t$, $B$, and presence of holes. Apart from the decrease in scaling exponent, we notice that enhancement vanishes for higher $B$ values as $W_{B}$ is too large (Figs. 4(b) and (c)). The many-body non-interacting probes are found to be more robust against disorder, as shown in Fig. 5 for the same four points in the $B-\alpha_{z}$ plane. Here we notice higher values of scaling exponent as well as enhancement even for higher $W_{B}$ values (Figs. 5(b), (c)). Therefore our analysis shows that even in the presence of imperfections, the sensing performance can show reduced advantage, which is encouraging from an experimental point of view.
Figure 4: Disorder for single-particle probes. $(B,\alpha_{z})=(0.05,0.1),(0.9,0.1),(0.9,0.9),(0.05,0.9)$ in (a), (b), (c), and (d), respectively. The red, blue, and green circles represent disorder in tunneling, magnetic field, and onsite potential, respectively.
Figure 5: Disorder for many-body probes. $(B,\alpha_{z})=(0.05,0.1),(0.9,0.1),(0.9,0.9),(0.05,0.9)$ in (a), (b), (c), and (d), respectively. The red, blue, and green circles represent disorder in tunneling, magnetic field, and onsite potential, respectively.
V Multi-parameter sensing
@@
Figure 6: Multi-parameter sensing. Probe state is the single particle ground state of the Hamiltonian in Eq. (7) with t=1t{=}1, and B=0.01tB{=}0.01t. The QFIM components are computed using Eq. (5). (a) The $(\alpha_{z},\alpha_{z})$ component of the QFIM $\mathcal{F}^{Q}$ across a range of $\alpha_{y}$ and $\alpha_{z}$ for L=100L{=}100. Four points chosen in the parameter space for scaling analysis are $(\alpha_{z},\alpha_{y})=(0.05,0.1),(0.9,0.1),(0.9,0.9),(0.05,0.9)$, denoted by the cross, diamond, plus, and circle, respectively. (b) Scaling of $(\mathcal{F}^{Q})_{\alpha_{z},\alpha_{z}}$ at the four points. (c) The $(\alpha_{y},\alpha_{y})$ component of $\mathcal{F}^{Q}$. (d) Scaling of $(\mathcal{F}^{Q})_{\alpha_{y},\alpha_{y}}$ at the four points. (e) The $(\alpha_{y},\alpha_{z})$ component of $\mathcal{F}^{Q}$. (d) Scaling of $(\mathcal{F}^{Q})_{\alpha_{y},\alpha_{z}}$ at the four points.
@@
In this section, we consider a more general scenario where we do not presume prior knowledge of the SOC parameters i.e. the Rashba fields in any direction. The goal is to estimate both $\alpha_{y}$ and $\alpha_{z}$ simultaneously, making the problem multi-parametric. To analyze the system’s sensitivity to changes in these parameters, we compute the multi-parameter QFIM $\mathcal{F}^{Q}$, which provides a comprehensive description of the ultimate precision limits for estimating multiple parameters concurrently. In Fig. 6(a) we shows the component $\mathcal{F}^{Q}_{\alpha_{z},\alpha_{z}}$ as a function of the two parameters. Fig. 6(b) shows that $\mathcal{F}^{Q}_{\alpha_{z},\alpha_{z}}$ reaches super-linear scaling in a wide range of $\alpha_{z}$ and $\alpha_{y}$ and can also attain close to Heisenberg scaling. Similar observation also holds for the component $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{y}}$, as shown in Figs. 6(c) and (d). From this analysis we can also infer that the quantum enhancement of the sensing of $\alpha_{y}$ when $\alpha_{z}$ is known, can reach the Heisenberg limit as well. In general, however, the sensitivity depends on the off-diagonal component $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{z}}$ that captures the correlation between the estimates of $\alpha_{z}$ and $\alpha_{y}$. Figs. 6(e) and (f) show that the scaling of $\mathcal{F}^{Q}_{\alpha_{y},\alpha_{z}}$ also stays close to quadratic across the whole range. In the following, we address how these QFI bounds can be obtained with a practical measurement basis.
VI Optimal measurement